OpenStudy (anonymous):

Your friend runs up to you, scared that he is not ready for the upcoming quadratics test. To help him study, you will create four different quadratic functions. Then demonstrate to him how to rewrite each function as a group of factors, if possible. The function f(x) is a difference of squares. The function g(x) is a sum of squares. The function h(x) is a perfect square trinomial. The function j(x) can only have a GCF factored out of it.

OpenStudy (anonymous):

\[f(x) = x^2-a^2 = (x+a)(x-a)\]

OpenStudy (anonymous):

\[g(x) = x^2+a^2\] this can be simplified into factors by using the quadratic formula but the factors would be complex numbers ( or imaginary numbers)

OpenStudy (anonymous):

\[h(x) = x^2 + 2ax +a^2\] this can be simplified as \[h(x) = (x+a)^2\] or \[h(x) = (x+a) (x+a)\]

OpenStudy (anonymous):

\[j(x)=ax^2+mx\]\[j(x)= x(ax+m) \]

OpenStudy (anonymous):

in j(x) the GCF is x

OpenStudy (anonymous):

are the options clear?

OpenStudy (anonymous):

yes so those would be my answers?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

factors are not possible for g(x)

OpenStudy (anonymous):

thaaaank you!!

OpenStudy (anonymous):

Its been a pleasure:)